Chapter 26, Problem 012 Near Earth, the density of protons in the solar wind (a stream of particles from the Sun) is 12.1 cm-3, and their speed is 559 km/s. (a) Find the current density of these protons. (b) If Earths Near Earth, the density of protons in the solar wind (a stream of particles from the Sun) is 12.1 cm , and their speed is 559 km/s. (a) Find the current density of these protons. (b) If Earths magnetic field did not deflect the protons, what total current Earth receive? (a) Number (b) Number Units SE UnitsNear Earth, the density of protons in the solar wind (a stream of particles from the Sun) is 12.1 cm-3, and their speed is 559 km/s. (a) Find the current density of these protons. (b) If Earth's magnetic field did not deflect the protons, what total current Earth receive?
Answer #1

A.

Current density is given by:

J = q*d*V

q = 1.6*10^-19

d = density of proton = 12.1 protons/cm^3 = 12.1*10^6 protons/m^3

V = 559 km/sec = 559000 m/sec

So,

J = 1.6*10^-19*12.1*10^6*559000

J = 1.082*10^-6 A/m^2

B.

J = i/A

i = J*A

A = pi*Re^2

i = 1.082*10^-6*pi*(6.37*10^6)^2

i = 1.379*10^8 Amp.